# Self-embedding similitudes of Bedford-McMullen carpets with dependent ratios

### v1

@@ 4. The dependent self-affine case
By Theorem 1.3, it suffices to consider the dependent case. Let $K,n,m$ be as at the beginning of this section. If $f\in\mathcal{S}(\mathbb{R}^{2})$ sends $K$ into itself, then the above theorem tells us that $f$ is not oblique. So it is not hard to find some integer $k\geq 1$ such that $f^{k}(x)=\lambda^{k}x+\eta$ for some $\eta\in\mathbb{R}^{2}$, i.e., the orthogonal part of $f^{k}$ is simply the identity matrix.
Note that $f^{k}$ sends the $x$-axis to a line with a unit vector $v=(1,0)$. If $N\geq 2$ but $\frac{\log\lambda}{\log n}\notin\mathbb{Q}$, we can apply Lemma 4.1 to $f^{k}$ and get (recall the second paragraph in Remark 4.2) some $g\in\mathcal{S}(\mathbb{R}^{2})$ of which the contraction ratio $\rho$ is an irrational power of $n$. However, applying the argument in the proof of the above theorem, we still arrive at (4.9) (where now $v=(1,0)$) and hence $\frac{\log\rho}{\log n}\in\mathbb{Q}$. This is a contradiction.
If $N=1$, then every row contains at most one selected rectangle in the initial patter of $K$. In particular, #$\#\Lambda=\#J\leq m<n$. So there must be a vacant column. On the other hand, since #$\#\Lambda\geq 2$ and $K$ is not contained in any vertical line, #$\#I\geq 2$. Therefore, $\pi_{1}(K)=E(n,I)$ is a deleted-digit set with Hausdorff dimension $10<\frac{\log\#I}{\log n}\leq\frac{\log\#\Lambda}{\log n}<$ 1. Noting that $f^{t}(K)\subset K$ implies that
$\lambda^{t}(E(n,I))+\pi_{1}(\eta)=\lambda^{t}(\pi_{1}(K))+\pi_{1}(\eta)=\pi_{1}(f^{t}(K))\subset\pi_{1}(K)=E(n,I).$
@@ 5. The dependent self-similar case
$=\begin{pmatrix}\frac{3}{5}&-\frac{4}{5}\\ -\frac{4}{5}&-\frac{3}{5}\end{pmatrix}\cdot K+(\tfrac{9}{5},\tfrac{18}{5})$
@@
    either $f\varphi_{R^{\prime}}(\ell_{t^{\prime}})\subset\varphi_{R}(\ell_{t})$ or $f\varphi_{R^{\prime}}(\ell_{t^{\prime}})\subset\varphi_{R}(\ell_{t+1})$.
@@
If (2) is false, we see by (1) that $\varphi_{R}(\ell_{t})$ meets $f\varphi_{R^{\prime}}(P)$ only at the vertex $\varphi_{R}(v_{t})$. So there is another level-$k^{\prime}$ square $R^{\prime\prime}$ such that $\varphi_{R}(v_{t})\in f\varphi_{R^{\prime\prime}}(P)$. By Corollary 5.3, there is a sequence $\{x_{p}\}_{p=1}^{\infty}\subset\ell_{t}$ such that $x_{p}\to v_{t}$ as $p\to\infty$. So $\varphi_{R}(x_{p})\to\varphi_{R}(v_{t})\in f\varphi_{R^{\prime}}(K)$. Since $\varphi_{R}(\ell_{t})$ meets $f\varphi_{R^{\prime}}(P)$ only at the vertex $\varphi_{R}(v_{t})$, $\varphi_{R}(x_{p})\notin f\varphi_{R^{\prime}}(K)$. But recalling that $\varphi_{R}(K)\subset f(K)$, we have for all large $p$ that $\varphi_{R}(x_{p})\in f(K)$. This implies that
@@
It remains to prove the existence of $R$. By Lemma 5.5, there is some $k_{0}$ and a level-$k_{0}$ square $Q$ such that $\varphi_{Q}(K)\subset f(K)$. Since $K$ is not contained in any line, we have #$\#I\geq 2$ and #$\#J\geq 2$ (here we abuse slightly these notations from Section 2.1). Iterating the initial pattern if necessary, we can find $(i_{1},j_{1}),(i_{2},j_{2}),(i_{3},j_{3})\in\Lambda$ such that
@@
If there is some $11\leq i\leq n-$ 1 such that $(i,\underline{j})\in\Lambda$ but $(i-1,\underline{j})\notin\Lambda$, then it suffices to take $R=\varphi_{Q}\varphi_{i_{1},j_{1}}\varphi_{i,\underline{j}}([0,1]^{2})$. If such a digit does not exist, then $(0,\underline{j})\in\Lambda$. If $I_{\underline{j}}\supsetneq\{0\}$, then $12\leq\#I_{\underline{j}}\leq n-$ 1, where the upper bound holds because $K$ does not contain horizontal segments. So there are at least $4$ level-$2$ squares in the bottom row and it is not hard to see that at least one of them, say $Q^{\prime}$, satisfies that the square of side $n^{-2}$ left adjacent to $Q^{\prime}$ is not selected. Then it suffices to take $R=\varphi_{Q}\varphi_{i_{1},j_{1}}(Q^{\prime})$. Similarly (looking at $(i_{2},j_{2})$ instead), it suffices to consider when $J_{\underline{i}}=\{0\}$.
@@
Let $f(x)=\lambda R_{\theta}x+a$ be as in Theorem 1.6 and let $K$ be as at the beginning of this section. If $\frac{\theta}{\pi}\in\mathbb{Q}$ and $f$ is an oblique self-embedding similitude of $K$, then $|\tan\theta|=1$.
@@
By Lemma 5.7 and the obliqueness of $f$, it suffices to show that $\tan\theta\in\mathbb{Q}$. Fix any $1\leq t\leq p$ and pick $R,R^{\prime},k,k^{\prime},t^{\prime}$ accordingly as in Proposition 5.6. Without loss of generality, assume that $f\varphi_{R^{\prime}}(\ell_{t^{\prime}})\subset\varphi_{R}(\ell_{t})$, which is the first possibility in Proposition 5.6(2).
@@
Since $\varphi_{R}(\ell_{t})$ is parallel to $\varphi_{R^{\prime}}(\ell_{t})$ and $\varphi_{R^{\prime}},\varphi_{R}$ involve no rotations, $f$ sends a line parallel to $\ell_{t^{\prime}}$ to another line parallel to $\ell_{t}$. Without loss of generality, assume that $t^{\prime}<t$. Then a simple geometric observation (see Figure 6) tells us that $R_{\theta}$ can be realized as follows: first rotate the line containing $\ell_{t^{\prime}}$ to the one containing $\ell_{t^{\prime}+1}$, and then to the one containing $\ell_{t^{\prime}+2}$ and so on until to $\ell_{t}$. Thus
@@
Again, we may assume that $K$ is not supported in any line. By Lemma 5.5, $\frac{\theta}{\pi}\in\mathbb{Q}$. Then the statement follows directly from the corollary. ∎

### v2

@@ 4. The dependent self-affine case
Let $K,n,m$ be as at the beginning of this section. If $f\in\mathcal{S}(\mathbb{R}^{2})$ sends $K$ into itself, then the above theorem tells us that $f$ is not oblique. So it is not hard to find some integer $k\geq 1$ such that $f^{k}(x)=\lambda^{k}x+\eta$ for some $\eta\in\mathbb{R}^{2}$, i.e., the orthogonal part of $f^{k}$ is simply the identity matrix.
Note that $f^{k}$ sends the $x$-axis to a horizontal line. If $\frac{\log\lambda}{\log n}\notin\mathbb{Q}$, we can apply Lemma 4.1 to $f^{k}$ and get (recall the second paragraph in Remark 4.2) some $g\in\mathcal{S}(\mathbb{R}^{2})$ of which the contraction ratio $\rho$ is an irrational power of $n$. However, if $N\geq 2$, applying the argument in the proof of the above theorem (where now $v=(1,0)$) , we can still arrive at (4.10) and hence $\frac{\log\rho}{\log n}\in\mathbb{Q}$. This is a contradiction.
If $N=1$, then every row contains at most one selected rectangle in the initial pattern of $K$. In particular, #$\#\Lambda=\#J\leq m<n$. So there must be a vacant column. On the other hand, since #$\#\Lambda\geq 2$ and $K$ is not contained in any vertical line, #$\#I\geq 2$. Therefore, $\pi_{1}(K)=E(n,I)$ is a deleted-digit set with Hausdorff dimension $10<\frac{\log\#I}{\log n}\leq\frac{\log\#\Lambda}{\log n}<$ 1. Noting that $f^{k}(K)\subset K$ implies that
$\lambda^{k}(E(n,I))+\pi_{1}(\eta)=\lambda^{k}(\pi_{1}(K))+\pi_{1}(\eta)=\pi_{1}(f^{k}(K))\subset\pi_{1}(K)=E(n,I).$
@@ 5. The dependent self-similar case
    either $f\varphi_{R^{\prime}}(\ell_{t^{\prime}})\subset\varphi_{R}(\ell_{t})$ or $f\varphi_{R^{\prime}}(\ell_{t^{\prime}+1})\subset\varphi_{R}(\ell_{t})$.
@@
If (2) is false, we see by (1) that $\varphi_{R}(\ell_{t})$ meets $f\varphi_{R^{\prime}}(P)$ only at the vertex $\varphi_{R}(v_{t})$. By Corollary 5.3, there is a sequence $\{x_{p}\}_{p=1}^{\infty}\subset\ell_{t}\cap K$ such that $x_{p}\to v_{t}$ as $p\to\infty$. So $\varphi_{R}(x_{p})\to\varphi_{R}(v_{t})\in f\varphi_{R^{\prime}}(K)$ and $\varphi_{R}(x_{p})\in\varphi_{R}(\ell_{t})$. Since $\varphi_{R}(\ell_{t})$ meets $f\varphi_{R^{\prime}}(P)$ only at the vertex $\varphi_{R}(v_{t})$, $\varphi_{R}(x_{p})\notin f\varphi_{R^{\prime}}(K)$. But recalling that $\varphi_{R}(K)\subset f(K)$, we have $\varphi_{R}(x_{p})\in f(K)$ for all $p$. This implies that
@@
It remains to prove the existence of $R$. By Lemma 5.5, there is some $k_{0}$ and a level-$k_{0}$ square $Q$ such that $\varphi_{Q}(K)\subset f(K)$. Since $K$ is not contained in any line, we have #$\#I\geq 2$ and #$\#J\geq 2$ (here we abuse slightly these notations from Section 2.1). Iterating the initial pattern if necessary, we may assume that there are $(i_{1},j_{1}),(i_{2},j_{2}),(i_{3},j_{3})\in\Lambda$ such that
@@
If there is some $11\leq i\leq n-$ 1 such that $(i,\underline{j})\in\Lambda$ but $(i-1,\underline{j})\notin\Lambda$, then it suffices to take $R=\varphi_{Q}\varphi_{i_{1},j_{1}}\varphi_{i,\underline{j}}([0,1]^{2})$. If such a digit does not exist, then $(0,\underline{j})\in\Lambda$. If $I_{\underline{j}}\supsetneq\{0\}$, then $12\leq\#I_{\underline{j}}\leq n-$ 1, where the upper bound holds because $K$ does not contain horizontal segments. So there are at least $4$ level-$2$ squares in the bottom row and it is not hard to see that at least one of them, say $Q^{\prime}$, satisfies that the grid square of side $n^{-2}$ left adjacent to $Q^{\prime}$ is contained in $[0,1]^{2}$ but unselected. Then it suffices to take $R=\varphi_{Q}\varphi_{i_{1},j_{1}}(Q^{\prime})$. So we may assume that $I_{\underline{j}}=\{0\}$. Similarly (looking at $(i_{2},j_{2})$ instead), it suffices to consider when $J_{\underline{i}}=\{0\}$.
@@
Let $f(x)=\lambda R_{\theta}x+a$ be as in Theorem 1.6. If $\frac{\theta}{\pi}\in\mathbb{Q}$ and $f$ is an oblique self-embedding similitude of $K$, then $|\tan\theta|=1$.
@@
By Lemma 5.7 and the obliqueness of $f$, it suffices to show that $\tan\theta\in\mathbb{Q}$. Fix any $1\leq t\leq p$ and pick $R,R^{\prime},k,k^{\prime},t^{\prime}$ accordingly as in Proposition 5.6. Without loss of generality, assume that $f\varphi_{R^{\prime}}(\ell_{t^{\prime}})\subset\varphi_{R}(\ell_{t})$, which is the first case in Proposition 5.6(2).
@@
Since $\varphi_{R}(\ell_{t})$ is parallel to $\varphi_{R^{\prime}}(\ell_{t})$ and $\varphi_{R^{\prime}},\varphi_{R}$ involve no rotations nor reflections, $f$ sends a line parallel to $\ell_{t^{\prime}}$ to another line parallel to $\ell_{t}$. Without loss of generality, assume that $t^{\prime}<t$ and $v_{1},\ldots,v_{p}$ are in counterclockwise order. Then a simple geometric observation (see Figure 7) tells us that $R_{\theta}$ can be realized as follows: first rotate the line containing $\ell_{t^{\prime}}$ to the one containing $\ell_{t^{\prime}+1}$, and then to the one containing $\ell_{t^{\prime}+2}$ and so on until to $\ell_{t}$. Thus
@@
Again, we may assume that $K$ is not supported in any line. By Lemma 5.5, $\frac{\theta}{\pi}\in\mathbb{Q}$. Then the statement follows directly from the above corollary. ∎
Acknowledgements. I thank Professor Huo-Jun Ruan for reading the manuscript carefully and pointing out several mistakes and a number of typos in an early version of it.